OpenStudy (anonymous):

Who wants medals! 2x + 3y = 3 3x – 2y = 11 What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

OpenStudy (whpalmer4):

Why don't you multiply the first equation on both sides by 3, and the second equation on both sides by 2, and add them together? You should find that you now have an equation with only 1 variable, which you can solve. Use that value in either of the original equations to find the other one. Then select the correct answer, and I'll give you a medal!

OpenStudy (whpalmer4):

Sorry, multiply the first equation by 2, the second equation by 3...

OpenStudy (amoodarya):

then sum of them deleted y so you solve for x

OpenStudy (anonymous):

is it x + 3y =6

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

No. \[2x + 3y = 3\]\[3x-2y = 11\] Multiply the first one by 2, and the second by 3: \[2*2x + 2*3y = 2*3\]\[3*3x - 3*2y = 3*11\] --------------------- 13x = 39 solve that for x, then use it to find y.